OpenStudy (anonymous):

Solve the equation: |2x − 5| less than or equal too 9

OpenStudy (anonymous):

can anyone figue this out

OpenStudy (anonymous):

-9<=2x-5<=9 -9+5<=2x<=9+5 -4<=2x<14 -2<=x<=7

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome dear